An individual has $30,000 invested in a stock with a beta of 0.6 and another $30,000 invested in a stock with a beta of 1.8. If these are the only two investments in her portfolio, what is her portfolio's beta? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Total investment=(30,000+30,000)=$60,000

Portfolio beta=Respective beta*Respective investment weight

=(30,000/60,000*0.6)+(30,000/60,000*1.8)

=1.2